# Unlocking Text Capabilities in Vision Models

### v2

@@ 5 Experiments
We first provide results on our newly formulated visual classifiers. We use the challenging ImageNet-1K dataset due to the widespread publicly available visual classifiers trained and evaluated on it. We apply our method on a diverse set of 40 visual classifiers. For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) [21], Wide ResNets [22], ResNeXts [23], ShuffleNetv2 [24], EfficientNetv2 [25], Densely Connected Networks (DenseNets) [26], ConvNeXts [27] and ConvNeXtv2 [28]. For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) [29], DINOv2 [30], BeiT [31], the hybrid Convolution-Vision Transformer CvT [32], Swin Transformer [33] and Swin Transformer v2 [34]. All models are pretrained on ImageNet-1K from the PyTorch [35] and HuggingFace [36] libraries. Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K. Models with a subscript v2 are trained following the updated PyTorch training recipe [37]. Finally, BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k. Both the pretrained classifier and text encoder remain frozen, only the MLP is trained on the ImageNet training set following Eq. 1.
Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K. For the text encoder, we use the MiniLM Sentence Encoder [38], with ablation studies on other text encoders presented in Section D of the Appendix. Results are presented in Table 1. We report the replicated Top-1 accuracy of the re-formulated classifier with our method in the first column, the original Top-1 accuracy of the classifier in the second column, and the difference between them ($\Delta$) in the last column. As can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models. We also perform ablation studies on the MLP to verify its effectiveness in Section C of the Appendix. We refer to Appendix Section F for results on other models and to Section G for implementation details.
@@
DenseNet169 75.46 75.60 --0.14 Swin-Base 83.22 83.58 --0.36
EfficientNetv2-S 84.04 84.23 --0.19 Swinv2-Base 83.72 84.11 --0.39
EfficientNetv2-M 84.95 85.11 --0.16 BeiT-B/16 84.54 85.06 --0.52
@@
ConvNeXt-Small 83.42 83.62 --0.20 DINOv2-B 84.40 84.22 +0.18
ConvNeXt-Base 83.88 84.06 --0.18 ConvNeXtV2-B 84.56 84.73 --0.17
@@
Table 1: Performance comparison of our re-formulated classifiers for several models. Top-1 indicates our results of the new formulation, and Orig. denotes the original Top-1 accuracy. $\Delta$ represents the difference between them ($\Delta$ = Top-1 -- Orig).
Zero-Shot Concept Bottleneck Models: We report CBM evaluation results on the ImageNet validation set using the top-1 accuracy in Table 2. As baselines, we compare with 9 methods: LF-CBMs [15], LaBo [16], CDM [17], DCLIP [39], DN-CBMs [18], DCBM [40], Multi-Modal (MM) CBMs [41], CF-CBM [42] and SALF-CBM [43]. All these methods are supervised and use CLIP-based models either for computing the concept activations or as a backbone. Our zero-shot CBMs (ZS-CBMs) outperforms all the supervised CBMs and sets a new state-of-the-art. Notably, even a simple ResNet-50 classifier trained solely on ImageNet already outperforms the CBM for the significantly more powerful ResNet-50 CLIP model trained on 400M samples. The best results are obtained by the ConvNeXtv2 model, which achieve a top-1 accuracy of $86.4$. All models show close to original accuracy, which means we can transform any classifier to be inherently interpretable without much loss in performance.
Figure 4: Qualitative examples of our zero-shot CBMs. We show the top-detected concepts, each with their corresponding importance score to the prediction.
Effectivenss of Concept Interventions: As a supplementary experiment, we also report concept intervention results on our ZS-CBMs in Section E of the Appendix, showing the effectiveness of the concepts, how we can mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
@@
LF-CBM CLIP ResNet50 67.5 DN-CBM CLIP ViT-B/16 79.5
LF-CBM CLIP ViT-B/16 75.4 SALF-CBM-S CLIP ViT-B/16 75.3
LaBo CLIP ResNet50 68.9 SALF-CBM-D CLIP ViT-B/16 76.3
LaBo CLIP ViT-B/16 78.9 DCBM-SAM2 CLIP ViT-L/14 77.9
CDM CLIP ResNet50 72.2 DCBM-GDINO CLIP ViT-L/14 77.4
CDM CLIP ViT-B/16 79.3 DCBM-RCNN CLIP ViT-L/14 77.8
DCLIP CLIP ViT-B/32 63.0 MM-CBM CLIP ViT-B/16 77.9
DCLIP CLIP ViT-B/16 68.0 MM-ProtoSim CLIP ViT-B/16 78.8
DCLIP CLIP ViT-L/14 75.0 CF-CBM-H CLIP ViT-B/16 77.4
DN-CBM CLIP ResNet50 72.9 CF-CBM-L CLIP ViT-B/16 78.4
@@
Table 2: Performance of Supervised and Zero-Shot CBMs on the ImageNet validation set.
@@
In Figure 4, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance. The weight importance is calculated by multiplying the concept activation with its corresponding weight to the predicted class. We use various concept sets to demonstrate the flexibility of our method to any desired concept set directly at test time (on-the-fly), as this process simply involves encoding the chosen concept set using the text encoder. All examples use the LF-CBM concept set [15]. By observing the first example, the image is predicted as a “scorpion" because it is has a lizard-like overall body, it is a venomous, desert animal with claws on its legs, and has large claws on its hand that look like a crab. Interestingly, in the second example, the top-detected concept is a “lift arm on the side.” Although this is not the primary feature defining a dumbbell, it reflects a well-documented bias in the literature [44] regarding the “dumbbell” class. Because most training images for this class show a dumbbell being lifted by an arm, the classifier not only learns to recognize the dumbbell but also associates it with the hand or arm that lifts it. With our method, we can obtain a textual interpretation of the biases that the original classifier learns. We also provide qualitative examples of global class-wise concepts detected in Section K of the appendix.
Zero-Shot Decoding of Visual Features: We now evaluate the performance of the textually decoded visual features. Since the ground-truth content of visual features is unknown—that is, we do not know exactly what the model uses for reasoning—we rely on COCO captions [45] for evaluation. While these captions may not perfectly correspond to the actual information represented by a visual feature vector—potentially differing from what a human annotator expects—they provide a reasonable approximation for this purpose, and represent the best available alternative for evaluation. Note that the COCO dataset differs in distribution than ImageNet, as a single image may contain many objects, interactions between them, and potentially categories not included in ImageNet (e.g., person). Therefore, it also serves as a way to evaluate generalization of our method to other datasets, given that we only used the ImageNet class names for training. Since we do not train any model on the ground-truth image captions provided by COCO, we use zero-shot image captioning as a benchmark.
We present results on the widely used “Karpathy test split" with various vision classifiers. As baselines, we compare our approach against existing methods in zero-shot image captioning, specifically ZeroCap [3] and ConZIC [46], both which use CLIP. For evaluation, we employ standard natural language generation metrics: BLEU-4 (B@4) [47], METEOR (M) [48], ROUGE-L (R-L) [49], CIDEr (C) [50], and SPICE (S) [51]. Results are shown in Table 3(a). ConvNeXtv2 achieves state-of-the-art performance on CIDEr and SPICE, the two most critical metrics for evaluating image captioning systems. Even with a simple ResNet-50 vision encoder trained on ImageNet-1K (1.2 million images), our approach outperforms the baseline methods on CIDEr and SPICE, despite the latter utilizing the significantly more powerful CLIP vision encoder, trained on 400 million image-text pairs. We also present qualitative examples of the decoded visual features for different models in Figure 5. From the first example, BeiT-L/16 captures features of both the vegetables and the dog, whereas ConvNeXtV2 captures only the vegetables. In contrast, ViT-B/16 and DINOv2 focus exclusively on the dog and its characteristics. Although some generations lacks semantic correctness (e.g., who was born in a dog), they are still meaningful enough for humans to understand and reason. This application allows human-friendly interpretations of visual features, understandable even to a layman user.
@@
(a) Zero-Shot Image Captioning Performance

### v3

@@ 4 Experiments
DenseNet169 75.46 75.60 -0.14 Swin-Base 83.22 83.58 -0.36
EfficientNetv2-S 84.04 84.23 -0.19 Swinv2-Base 83.72 84.11 -0.39
EfficientNetv2-M 84.95 85.11 -0.16 BeiT-B/16 84.54 85.06 -0.52
@@
ConvNeXt-Small 83.42 83.62 -0.20 DINOv2-B 84.40 84.22 +0.18
ConvNeXt-Base 83.88 84.06 -0.18 ConvNeXtV2-B 84.56 84.73 -0.17
@@
Table 1: TextUnlock-ed visual classifiers maintain classification performance. Comparison of our re-formulated classifiers for several models. Top-1 indicates our results of the new formulation, and Orig. denotes the original Top-1 accuracy. $\Delta$ represents their difference ($\Delta$ = Top-1 - Orig).
We first provide results of the the classifier with its corresponding vision-language distribution aligned using our method TextUnlock. We use the ImageNet-1K benchmark dataset due to the widespread publicly available visual classifiers trained and evaluated on it. In Section F of the Appendix, we also report results on other datasets including Places365 (zhou2017places), EuroSAT (helber2019eurosat) and DTD (Cimpoi2013DescribingTI) showing that our method is also applicable to domain-specific and fine-grained datasets. We apply TextUnlock on a diverse set of 40 visual classifiers. For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) (He2015DeepRL), Wide ResNets (Zagoruyko2016WRN), ResNeXts (Xie2016AggregatedRT), ShuffleNetv2 (Ma_2018_ECCV), EfficientNetv2 (Tan2021EfficientNetV2SM), Densely Connected Networks (DenseNets) (Huang2016DenselyCC), ConvNeXts (Liu2022ACF) and ConvNeXtv2 (Woo2023ConvNeXtVC). For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) (dosovitskiy2021an), DINOv2 (oquab2024dinov), BeiT (bao2022beit), the hybrid Convolution-Vision Transformer CvT (Wu2021CvTIC), Swin Transformer (Liu2021SwinTH) and Swin Transformer v2 (Liu2021SwinTV). All models are pretrained on ImageNet-1K from the PyTorch (torchvision2016) and HuggingFace (wolf-etal-2020-transformers) libraries. Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K. Models with a subscript v2 are trained following the updated PyTorch training recipe (PytorchNewTrainRecepie). Finally, BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k. When training with TextUnlock, both the pretrained classifier and text encoder remain frozen, only the MLP is trained on the ImageNet training set following Eq. 1.
Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K. For the text encoder, we use the MiniLM Sentence Encoder (wang2020minilm) as it is powerful, fast and efficient. We provide ablation studies on other text encoders in Section D of the Appendix. Results are presented in Table 1. We report the replicated Top-1 accuracy of the re-formulated classifier with our method in the first column, the original Top-1 accuracy of the classifier in the second column, and the difference between them ($\Delta$) in the last column. As it can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models. We also perform ablation studies on the MLP to verify its design, impact, role and that it learns meaningful transformations in Section C of the Appendix (see Section C.1 for the role and impact of the MLP and Section C.2 for its design hyperparameters). We also evaluate our transformed classifier robustness to prompt variations in Section H of the Appendix. For results on other models, we refer to Section L of the Appendix, and for implementation details, to Section J of the Appendix.
[…]
Effectivenss of Concept Interventions. Another common way to evaluate interpretability of CBMs is concept intervention. We report concept intervention results on our ZS-CBMs in Section E of the Appendix. In this way, we show the effectiveness of the concepts, how we can mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
@@
LF-CBM CLIP ResNet50 67.5 LF-CBM CLIP ViT-B/16 75.4
LaBo CLIP ResNet50 68.9 LaBo CLIP ViT-B/16 78.9
CDM CLIP ResNet50 72.2 CDM CLIP ViT-B/16 79.3
DCLIP CLIP ResNet50 59.6 DCLIP CLIP ViT-B/16 68.0
DN-CBM CLIP ResNet50 72.9 DN-CBM CLIP ViT-B/16 79.5
DCBM-SAM2 CLIP ViT-L/14 77.9 DCBM-RCNN CLIP ViT-L/14 77.8
@@
Table 2: Our zero-shot CBMs outperform CLIP-based counterparts. Accuracy of Supervised and Zero-Shot CBMs (ZS-CBMs, ours) on ImageNet validation set. Similar backbones are color-coded. Best within backbone family is underlined, overall best is bold.
@@
In Figure 3, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance on the x-axis. The weight importance is calculated by multiplying the concept activation with its corresponding weight to the predicted class. We use various concept sets to demonstrate the flexibility of our method to any desired concept set directly at test time (on-the-fly), as this process simply involves encoding the chosen concept set using the text encoder. All examples use the LF-CBM concept set (oikarinen2023labelfree). By observing the first example, the image is predicted as a “scorpion" because it is has a lizard-like overall body, it is a venomous, desert animal with claws on its legs, and has large claws on its hand that look like a crab. Interestingly, in the second example, the top-detected concept is a “lift arm on the side.” Although this is not the primary feature defining a dumbbell, it reflects a well-documented bias in the literature (Samek2019TowardsEA) regarding the “dumbbell” class. Because most training images for this class show a dumbbell being lifted by an arm, the classifier not only learns to recognize the dumbbell but also associates it with the hand or arm that lifts it. With our method, we can obtain a textual interpretation of the biases that the original classifier learns. We also provide qualitative examples of global class-wise concepts detected in Section P of the appendix.
Figure 3: Qualitative examples of our zero-shot CBMs. We show the top-detected concepts, each with their corresponding importance score to the on the x-axis.
5 Zero-Shot Image Captioning
We also show that TextUnlock enables zero-shot image captioning with any pretrained visual classifier beyond CLIP models. Existing zero-shot captioning methods largely rely on CLIP and its shared vision–language space. Thanks to TextUnlock, zero-shot image captioning can now be performed for any pretrained visual classifier. We adapt the method introduced in ZeroCap (Tewel2021ZeroCapZI) for our purpose. Specifically, ZeroCap is a test-time approach that learns to produce a caption that maximizes the similarity with the image features. We first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$. Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text. We utilize an off-the-shelf pretrained language decoder model (e.g., GPT-2), denoted as $G$, to generate open-ended text. We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning (Li2021PrefixTuningOC) to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$. Specifically, we attach a set of learnable “virtual" tokens to $G$. Denote the generated output text of $G$ for one iteration as $h^{j}$, where $j$ represents the iteration number. $h^{j}$ is encoded with the text encoder $T$, to yield a vector $y^{j}$, That is, $y^{j}=T(h^{j})$. As $y^{j}$ and $\tilde{f}$ are now in the same embedding space, we maximize the cosine similarity between them in order to update the learnable tokens. We perform this process for several iterations. As this process is not the core contribution of our work, we leave more details to Section M of the appendix, which also includes a detailed illustration in Figure 2.
[…]
Figure 4: Qualitative examples of zero-shot image captioning.
@@ 5 Zero-Shot Image Captioning
[…]
Table 3: Zero-Shot Image Captioning Performance
Note that our results in Table 3 are outperformed by the baseline ZeroCap on the BLEU-4 (B4) and METEOR (M) metrics. However, it is important to note that B4 and M are n-gram overlap-based metrics. They assume that the generated caption follows a specific structure and style. We verify this hypothesis by applying compositional image captioning (Kulkarni2011BabyTU; Lu2018NeuralBT) with in-context learning. Specifically, a set of image-grounded verbs and concepts (e.g., attributes, objects) are first detected using the concept discovery step described in Section 3.2, and they are fed along with six in-context examples from the COCO captioning set to a language model, capable of imitating the style and structure of the in-context examples, in order to compose them into a natural sounding sentence. This also demonstrates the direct application of our approach to compositional image captioning. The last row of Table 3 with the superscript com shows that the (B4, M and R-L) are boosted, which verifies our hypothesis about the low scores of B4 and M compared to baseline methods. We provide more details and results on other models on this paradigm of compositional image captioning in Section K of the Appendix. We further show how we can use this approach to generate captions tailored to any domain (see Section K.1 in the Appendix).
6 Conclusion
We introduced a method for transforming any frozen visual classification model into a Concept Bottleneck Model (CBM). We proposed TextUnlock, the core of our method, that aligns the distribution of the original classifier with that of its vision–language counterpart. This method further enables a CLIP-free, label-free conversion of the frozen visual classification model into a CBM, and further allows us to derive zero-shot concept bottleneck concept-to-class classifiers (without training the linear classifier), Applied on 40 different models, our method outperforms supervised CLIP-based CBMs and achieves state-of-the-art results. TextUnlock further allows us to perform zero-shot image captioning on any visual classification model, also achieving state-of-the-art results and surpassing zero-shot captioning baselines based on CLIP. Finally, as with any research work, this study has its own limitations which are discussed in Section B of the Appendix.
Appendix A Pseudocode of TextUnlock
We provide a PyTorch-like pseudocode of TextUnlock in Listing 1.
# mlp: trainable MLP from visual_dim -> text_dim
# images: batch of B images, shape (N, 3, height, width)
visual_feats = vision_encoder(images) # (N, visual_dim)
logits = visual_feats @ classifier # (N, num_classes)
original_dist = softmax(logits, dim=-1) # (N, num_classes)
mapped_feats = mlp(visual_feats) # (N, text_dim)
mapped_feats = l2_norm(mapped_feats) # (N, text_dim)
text_feats = l2_norm(text_feats) # (N, text_dim)
pred_logits = mapped_feats @ text_feats.T # (N, num_classes)
pred_dist = softmax(pred_logits, dim=-1) # (N, num_classes)
# cross entropy with original model’s soft distribution
loss = -(original_dist * log(pred_dist)).sum(dim=1).mean()
loss.backward() # only mapper parameters are updated
Listing 1: PyTorch-like pseudocode for TextUnlock
Appendix B Limitations
As with any research work, our study has its own limitations that should be transparent and acknowledged. In particular, we identify a primary limitation of our method concerning wrong semantic associations of class names in CBMs. This issue is closely tied to the choice of concept set used. When using the 20K most common words in English as our concept set, we observe that class names get associated to wrong semantically-related concepts. Figure 1 illustrates such cases. In the first example, the bird drake is linked to artist-related concepts (Rihanna, Robbie, lyric). This occurs because the bird “drake" is less familiar to the text encoder than the artist “drake". In fact, a google search with the word “drake” yields directly the artist rather than the bird. In the second example, the top-detected concepts for the prediction “african grey” are incorrect semantic associations with the word “african” (ethiopian, tanzania) and the word "grey" (purple, blue) that do not pertain to the bird itself. We also observe similar cases that may raise ethical concerns (e.g., the animal “cock" leads to associations with male reproductive terms). For each example, we also report the total logit score of the prediction.
However, it is worth noting the following:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Figure 1: Limitations of our method in wrong semantic concept association
Appendix C Ablation Studies of the MLP
C.1 Role and Impact of the MLP
We first evaluate the role and impact of the MLP to verify its contribution and that it learns meaningful transformations, and does not collapse into a trivial function. We perform the following ablation studies: 1) Mean Ablation: For an image, we replace the input features to the MLP with a constant mean of the features calculated across the full ImageNet validation set. 2) Random Features: For an image, we replace the input features to the MLP with random values sampled from a normal distribution with a mean and standard deviation equal to that of the features calculated across the full ImageNet validation set. 3) Random Weight Ablation: We randomize the weights of the MLP projection. 4) Shuffled Ablation: For an image, we replace the input features to the MLP with input features of another random image in the validation dataset.
ResNet101v2 Ours 81.49 81.49 81.49 81.49
Ablated 0.10 0.11 1.70 0.11
ConvNeXt-Base Ours 83.88 83.88 83.88 83.88
Ablated 0.10 0.11 1.79 0.10
BeiT-L/16 Ours 87.22 87.22 87.22 87.22
Ablated 0.10 0.11 1.87 0.11
DINOv2-B Ours 84.40 84.40 84.40 84.40
Ablated 0.10 0.13 1.76 0.09
Table 1: Ablation studies of the MLP.
C.2 MLP Design
Next, we perform an ablation study over the number of layers and output dimension scaling factor (dim_out_factor) of the MLP. dim_out_factor specifies how much to scale the previous layer’s dimensionality. For example, in a ViT-B/16 model the hidden dimension is 768; setting dim_out_factor = 2 therefore expands the projection from 768 to 2 $\times$ 768 = 1536. Results are shown in Table 2 for a ResNet50 model. We observe that the 2 layers and a output dimension scaling factor of 2 provides the best results.
Layers dim_out_factor Top-1 (%)
1 1 72.48
1 2 74.01
2 1 75.41
2 2 75.80
Table 2: Performance comparison of the MLP for different layer and dimension configurations.
Appendix D Ablation Studies on the Text Encoder
In Table 3, we present ablation studies using other text encoders from the Sentence-BERT library (reimers-2019-sentence-bert) with the ResNet50 visual classifier. We observe that the choice of the text encoder has minimal effect on the performance. This is because even lower-performing text encoders are capable of understanding class names.
Text Encoder Top-1 (%)
DistilRoberta 75.73
MPNet-Base 75.78
MPNet-Base-MultiQA 75.76
MiniLM 75.80
Table 3: Ablation studies on other text encoders
Appendix E Effectiveness of Concept Interventions
As a supplementary experiment, we also report concept intervention results on our ZS-CBMs. Interventions on CBMs are an effective tool to mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions. The Waterbirds-100 dataset (Sagawa2020Distributionally) is a standard dataset used in previous works (Rao2024DiscoverthenNameTC) to conduct CBM intervention experiments. It is a binary classification dataset of two classes: waterbirds and landbirds. The training images of waterbirds are on water backgrounds, and training images of landbirds are on land backgrounds. However, the validation images do not have that correlation, where waterbird images appear on land backgrounds and landbird images on water backgrounds. The model is assumed to learn the water–land background correlation to perform this classification task. By building a CBM, we can correct this bias by intervening in concepts in the CB layer. We curated a validation dataset of waterbirds/landbirds directly from the ImageNet validation set. Our curated dataset includes 140 validation images (70 for each class). We create our ZS-CBM using the two class prompts:“an image of a waterbird" (for the waterbird class) and “an image of a landbird" (for the landbird class), and using the same concept set from (Rao2024DiscoverthenNameTC) which includes a collection of bird-related concepts and a collection of land-related concepts. The ZS-CBM achieves a low accuracy, as shown in Table 4, which indicates the water-land bias the model performs for classification. To correct this, we intervene in the concepts in the CB layer, following the setup from (Rao2024DiscoverthenNameTC). Intervention R (Int. R): We zero-out activations of any bird concepts from the bottleneck layer, and expect the accuracy to drop. Intervention K (Int. K): We keep activations of bird concepts as they are, but scale down the activations of all remaining concepts by multiplying them with a factor of 0.1, and expect the accuracy to increase. Results are presented in Table 4 for some models, and demonstrate the success of our intervention experiments.
We also conduct concept intervention experiments with a more challenging multi-class setup. Unlike the Waterbirds dataset where bias‑correlation issues are assumed, we make no such assumption here. We instead evaluate whether intervening on class-related concepts in the bottleneck layer and zeroing out their activations, impairs model accuracy. A drop in accuracy indicates that these concepts are important for prediction. We select a subset of 10 classes from ImageNet following (Howard_Imagenette_2019). We use the original ImageNet validation images for those classes, rather than the validation set from (Howard_Imagenette_2019), because the latter contains original ImageNet training examples that our model has already seen. We employ an LLM to generate 5 highly-relevant concepts for each class, achieving in total 50 concepts. The classes and concepts we used are provided in Section Q of the supplementary material. We then measure CBM accuracy before and after intervention. Results are provided in Table 5. For each image, we zero out the activations of its class-related concepts and report the Intervention R (Int. R) metric. As shown in Table 5, this intervention reduces accuracy by approximately 20% on average, underscoring the concepts importance to the CBM.
[…]
BeiT-B/16 54.29 41.43 (-12.86) 58.57 (+4.28)
[…]
DiNOv2 52.86 43.57 (-9.29) 59.29 (+6.43)
BeiT-L/16 52.86 44.29 (-8.57) 58.57 (+5.71)
Table 4: CBM Interventions on the standard Waterbirds dataset
[…]
ResNet50 96.80 76.00 (-20.8)
ResNet101 97.00 76.80 (-20.2)
DINOv2-B 98.60 78.40 (-20.2)
BeiT-B/16 98.60 78.80 (-19.8)
[…]
Table 5: CBM Interventions in a multi-class setup
Appendix F Transformed Classifier Results on Other Datasets
We conduct extra experiments on 3 datasets. Places365 (domain-specific to scenes), DTD (domain-specific to texture and fine-grained), and EuroSAT (domain-specific to satellite images). For each dataset, we show the top-1 accuracy of the classifier’s original performance and the top-1 accuracy of our transformed classifier. Results are shown in Table 6.
Dataset Model Original (%) Transformed (%)
Places365 ResNet50 54.77 53.90
DenseNet161 56.13 55.54
EuroSAT ResNet50 93.95 94.23
WideResNet101 93.78 93.88
ViT-B/16 93.67 93.53
DTD ResNet50 69.47 69.26
WideResNet101 68.62 68.14
ViT-B/16 69.95 69.68
Table 6: Results of our transformed classifier on Places365, EuroSAT, and DTD datasets
Appendix G CBM Results on Other Datasets
We report CBM results on additional datasets: Places365 (domain-specific to scenes), DTD (domain-specific to texture and fine-grained), and EuroSAT (domain-specific to satellite images). When a baseline method is also reported on that dataset, we include it as a baseline. Otherwise, we use CLIP models as a baseline to act as a feature extractor and to compute concept activations, and train a linear classifier on top of the concept activations, formulating a CLIP supervised baseline. All baselines use the same concept set. Results are shown in Table 7. For Places 365, we can see that the transformed DenseNet161 classifier trained only on ImageNet, outperforms supervised CLIP-based ResNet and ViTs CBM methods in a zero-shot manner. The same applies for EuroSAT and DTD with our transformed Image-Net only trained classifiers. Therefore, the experiments show that our method scales to domain-specific and fine-grained datasets.
Dataset Method Model Acc (%)
Places365 Supervised
DCLIP CLIP-ResNet50 37.90
DCLIP CLIP-ViT-B/16 40.30
LF-CBM CLIP-ResNet50 49.00
LF-CBM CLIP-ViT-B/16 50.60
CDM CLIP-ResNet50 52.70
CDM CLIP-ViT-B/16 52.60
Zero-Shot
Ours ResNet50 51.57
Ours DenseNet161 53.42
EuroSAT Supervised
Baseline CLIP-ResNet50 86.27
Baseline CLIP-ViT-B/16 88.57
Zero-Shot
Ours ResNet50 94.22
Ours WideResNet101 94.12
Ours ViT-B/16 93.65
DTD Supervised
Baseline CLIP-ResNet50 57.77
Baseline CLIP-ViT-B/16 61.86
Zero-Shot
Ours ResNet50 68.88
Ours WideResNet101 66.97
Ours ViT-B/16 68.46
Table 7: CBM results on Places365, EuroSAT, and DTD datasets
Appendix H Prompt Variations
We evaluate the robustness of our transformed classifier to variations in text prompts. Using the transformed ViT-B/16, we provide a diverse set of prompts and measure the Top-1 accuracy on ImageNet. Results are shown in Table 8. We order the prompts from highest to lowest scoring. We find that our transformed classifier is robust to text prompts. The worst prompt degrades the baseline accuracy by only 0.36 points
Prompt Top-1 Accuracy (%)
an image of a {} 80.70
a photo of one {}. 80.67
a photo of a {}. 80.66
a close-up photo of a {}. 80.63
a black and white photo of a {} 80.63
a close-up photo of the {}. 80.62
a cropped photo of a {}. 80.61
a good photo of a {}. 80.61
a dark photo of a {}. 80.61
a bright photo of a {}. 80.60
a bright photo of the {}. 80.59
a bad photo of a {}. 80.57
a blurry photo of a {}. 80.57
a pixelated photo of a {}. 80.55
a photo of many {}. 80.54
a low-resolution photo of the {} 80.54
a low-resolution photo of a {}. 80.52
a photo of my {}. 80.45
a jpeg-corrupted photo of a {}. 80.34
Table 8: Effect of different text prompts on Top-1 accuracy using ViT-B/16.
Appendix I Other Concept Sets
Early works such as label-free CBMs use LLMs to automatically generate a set of concepts for a target class. In recent works such as DN-CBM (Rao2024DiscoverthenNameTC) and DCBM (Knab2024DCBMDV), it was however shown that general, pre-defined concept sets outperform the LLM-generated ones, which introduce many spurious correlations and hallucinated associations. In Table 9, we compare general pre-defined and LLM-generated concept sets. Across different types of architectures, general, pre-defined concept sets outperform the LLM-generated ones by several percent. Please note that while earlier works such as LF-CBMs employ LLM-generated concept sets (which underperform compared to the general, pre-defined concept set), all results reported in our main manuscript—including those from prior works—are presented using the general, pre-defined concept set to ensure a fair comparison.
Model Predefined LLM-generated
Swinv2-Base 82.60 80.34
ConvNext-Base 82.80 80.24
[…]
ViT-B/16 79.30 76.37
Table 9: Comparison of Predefined and LLM-generated concept sets on CBMs
Appendix J Implementation Details
For the text encoder, we use the all-MiniLM-L12-v1222https://huggingface.co/sentence-transformers/all-MiniLM-L12-v1 model available on the Sentence Transformers library (reimers-2019-sentence-bert). This text encoder was trained on a large and diverse dataset of over 1 billion training text pairs. It contains a dimensionality of $m=384$ and has a maximum sequence length of 256.
Our MLP projector is composed of 3 layers, the first projects the visual feature dimensions $n$ to $n\times 2$ and is followed by a Layer Normalization (Ba2016LayerN), a GELU activation function (Hendrycks2016GaussianEL) and Dropout (Srivastava2014DropoutAS) with a drop probability of 0.5. The second layer projects the $n\times 2$ dimensions to $n\times 2$ and is followed by a Layer Normalization and a GELU activation function. The final linear layer projects the $n\times 2$ dimensions to $m$ (the dimensions of the text encoder). We train the MLP projector with a batch size of 256 using the ADAM optimizer (KingBa15) with a learning rate of 1e-4 that decays using a cosine schedule (loshchilov2017sgdr) over the total number of epochs. We follow the original image sizes that the classifier was trained on.
Appendix K Compositional Image Captioning
Compositional Image Captioning is an old image captioning paradigm termed (Kulkarni2011BabyTU), later revived with deep learning methods (Lu2018NeuralBT). In compositional captioning, a set of image-grounded concepts (such as attributes, objects and verbs) are first detected, and a language model is then used to compose them into a natural sounding sentence. With the current advancements of Large Language Models (LLMs) and their powerful capabilities, we use an LLM as a compositioner. Specifically, we detect the top concepts and verbs to the image using the concept discovery method introduced in Section 3.2 and shown in Figure 2(a), and feed them, along with their similarity scores, to an LLM. For the concepts, we use the same concept set as in Section 3.2. We also add a list of the most common verbs (commonverbs) in English to the pool. This allows us to cover all possible words and interactions. We prompt the LLM to utilize the provided information to compose a sentence, given a limited set of in-context examples from the COCO captioning training set (in our experiments, we use 9 examples). This allows us to generate sentences adhering to a specific style and structure. For this experiment, we used GPT4o-mini (gpt4omini) as our LLM, as it is fast and cost-efficient. In the prompt, we explicitly instruct the LLM to refrain from reasoning or generating content based on its own knowledge or assumptions, and that all its outputs must be strictly grounded in the provided concepts, verbs, and score importance. We use the following prompt:
Results are shown in Table 10.